\begin{thebibliography}{AAA111}
\bibitem{Al} {\sc Albeverio. S.} {\it Theory of Dirichlet forms and applications,} pp. 3-106 in S. Albeverio, W. Schachermayer, M. Talagrand. Lectures on Probability theory and Statistics. St. Floor, 2000. Springer (2003). 
\bibitem{ABrMa} {\sc Albeverio. S, Brzezniak. Z and Mazzuchi. S.} {\it Probabilistic integrals: mathematical aspects. } Scholarpedia, 12(5):10429. (2017)
\bibitem{ACDP} {\sc Albeverio. S, Cordoni. F, Di Persio L. and Pellegrini. } {\it Asymptotic expansions for some locality models arising in finance. } Preprint (2017).
\bibitem{AlDiPMa} {\sc Albeverio. S, Di Persio L. and Mastrogiacomo. E.} {\it Small noise asymptotic expansions for
stochastic PDE's I. The case of a dissipative polynomially bounded
nonlinearity. } Tohoku. Math. J., {\bf  63} (2011), pp. 877--898.
\bibitem{ALDIBOUB} {\sc Albeverio. S, Dipersio. L, Mastrogiacomo. E and Smii. B.} {\it Explicit invariant measures for infinite dimensional SDE driven by L\'evy noise with dissipative nonlinear drift I.} Comm. Math. Sci.  Vol. 15, No. 4, pp. 957-983 (2017).
\bibitem{ALDIBOUB1} {\sc Albeverio. S, Dipersio. L, Mastrogiacomo. E and Smii. B.} {\it A class of L\'evy driven SDEs and their explicit invariant measures.} Pot. Anal. V. 45 (2), pp. 229-259, (2016).
\bibitem{AGoWu} {\sc  Albeverio. S, Gottschalk. H and J-L. Wu. } {\it Convoluted Generalized White noise, Schwinger Functions and their Analytic continuation to
 Wightman Functions.} Rev. Math. Phys, Vol. 8, No. 6, 763-817, (1996).
  \bibitem{AHk} {\sc Albeverio. S, Hilbert. A and Kolokoltsov. V.} {\it Uniform asymptotic bounds for the heat kernel and the trace of a stochastic geodesic flow.} Stochastics. 84(2-3): 315-333. (2012).
 \bibitem{AlGYo} {\sc Albeverio. S,  Gottschalk. H and Yoshida. M.W. } {\it Systems of classical particles in the Grand canonical ensemble, scaling limits and
 quantum field theory.}  Rev. Math. Phys, Vol. 17, No. 02, 175-226, (2005).
  \bibitem{AHKS} {\sc Albeverio. S,  Hoegh-Krohn.R and Streit.L. } {\it Energy forms, Hamiltonians and distorted Brownian paths.} J. Math.Phy. 18, 907 (1977).  
 \bibitem{AKKR} {\sc Albeverio. S, Kondratiev. Y, Kozitsky. Y and R\"{o}ckner. M.} {\it The statistical mechanics of quantum lattice systems. A path integral approach.} EMS (2009).
 \bibitem{ALiMa} {\sc Albeverio. S, Lytvynov. E and Mahnig.A.} {\it A model of the term structure of interest rates based on L\'evy fields." Stoch. Proc.
 Appl. 114.2 (2004): 251-263.}
 \bibitem{AlMa} {\sc Albeverio. S and Marinelli. C.} {\it Reconstructing the drift of a diffusion from partially observed transition probabilities.} Stoch. Proc. Appl. 115, pp. 1487-1502 (2005).
\bibitem{ALEBOU} {\sc Albeverio. S, Mastrogiacomo. E and Smii.
B.\,}{\it Small noise asymptotic expansions for stochastic PDE's
driven by dissipative nonlinearity and L\'evy noise.} Stoch. Proc.
Appl. 123(2013), 2084-2109.
%\bibitem{ABRM} {\sc Albeverio. S, Mandrekar. V and R$\ddot{u}$diger B.}
%{\it Existence of mild solutions for stochastic differential
%equations and semilinear equations with non-Gaussian L\'evy noise.}
%Stochastic Proc. Appl. 119 (2009), no. 3, 835--863.
\bibitem{ALRS} {\sc Albeverio. S, R\"{o}ckle. H and Steblovskaya. V.} {\it Asymptotic expansions for Ornstein-Uhlenbeck semigroups perturbed by potentials over Banach spaces.} Stoch and Stoch. Rep. 69, pp. 195-238, (2000).
\bibitem{ALBO} {\sc Albeverio. S and Smii. B.\,}{\it Asymptotic expansions for SDE's with small multiplicative
noise.} Stoch. Proc.
Appl. 125(2015), no.3, 1009-1031 .
\bibitem{ASte}{\sc Albeverio. S and Steblovskaya. V.} {\it A model of financial market with several interacting assets. Complete market case." Finance and Stochastics 6.3 (2002): 383-396.}
%\bibitem{AWZ}{\sc Albeverio. S ,  Wu. J-L and  Zhang. T. S.}
%{\it Parabolic SPDEs driven by Poisson white noise.} Stoch. Proc.
%Appl. 74 (1998), 21-36.
%\bibitem{DA} {\sc  Applebaum. D .} {\it L\'evy Processes and
%stochastic Calculus.} Cambridge University Press 2004.
\bibitem{Az} {\sc Azencott. R.} {\it Densit\'e  des diffusions en temps petit, d\'eveloppement asymptotique. } S\'em. Prob. XVIII, LN. Math. 1059, 402-494 (1984).
\bibitem{Ba} {\sc Baldi. P and Carmellino. l.} {\it Asymptotic of hitting probability for general one-dimensional pinned diffusions.} Ann. Appl. Prob. 12, 1071-1095 (2002). 
% \bibitem{BN} {\sc Barndorff. N. O.E and  Shephard. N.}{\it Non-Gaussian Ornstein-Uhlenbeck-based models and some of their uses in financial economics.}
% J. R. Stat. Soc. Ser. B Stat. Methodol. 63, no. 2, 167-241 (2001).
 %\bibitem{BB} {\sc Balian. R and Bloch. C.} {\it  Solutions of the Schr\"{o}dinger equation in terms of classical paths, Ann. of Phys. 85 (1974).  }
 \bibitem{BE} {\sc Ben Arous. G.} {\it M\'ethodes de Laplace et de la phase stationnaire sur l'espace de Wiener.} Stochastics. 25, 125-153 (1988).
 \bibitem{BBF} {\sc Berestycki. H. Busca. J. and Florent. I.} {\it Asymptotics and calibration of local volatility models.} Quant. Finan. 2, 61-69 (2002).
% \bibitem{Bi} {\sc Biagini. F.} {\it Stochastic calculus for fractional Brownian motion and applications. Springer, 2008.}
% \bibitem{BoLe} {\sc Boyarchenko. S and Levendorskii. S.Z.} {\it Non-Gaussian Merton-Black-Scholes Theory. Vol. 9. Singapore: World Scientific, 2002.}
%\bibitem{BRHA} {\sc Brze$\acute{z}$niak. Z and Hausenblas. E.} {\it
%Uniqueness in law of the It$\hat{o}$ integral with respect to L\'evy
%noise,} pp. 37-57 in Seminar Stoch. Anal., Random Fields and Appl.,
%VI, Birkhauser, Basel (2011).
\bibitem{CL}{\sc Capriotti. L.} {\it The Exponent Expansion: An effective approximation of transition probabilities of diffusion processes and pricing kernels of financial derivatives.} Int. J. Theor. Appl.Fin. (2006).
\bibitem{Co1}{\sc Colin Y. de Verdi\'ere} {\it Spectre du Laplacien et longueurs des g\'eod\'esiques p\'eriodiques.I, Composio Mathematica 27-1 (1973), 83-106.  }
\bibitem{Co2}{\sc Colin Y. de Verdi\'ere} {\it Spectre du Laplacien et longueurs des g\'eod\'esiques p\'eriodiques.II, Composio Mathematica 27-1 (1973), 83-106.  }
%\bibitem{CoTa}{\sc Cont. R and Tankov. P  }{\it Financial modelling with jump processes. }Vol. 2. CRC Press, 2004.
%\bibitem{DA}{\sc Da Prato. G and Zabczyk.J. }{\it Ergodicity for infinite-dimensional systems, London Mathematical Society Lecture Note
%Series},vol. 229, Cambridge University Press,   Cambridge,(1996).
%\bibitem{DGZ}{\sc Da Prato. G and Zabczyk. J. }{\it Ergodicity for infinite dimensional systems. Vol. 229.}Cambridge University Press, 1996.
\bibitem{Doss}{\sc Doss. H.} {\it Sur une r\'{e}solution stochastique de l'\'{e}quation de Schr\"{o}dinger \`{a} coefficients analytiques.} Comm. Math.Phys. V 73, 3, pp. 247-264 (1980).
%\bibitem{EG}{\sc Eberlein. E and Glau. K.  }{\it Variational Solutions of the Pricing PIDEs for European Options in L\'evy Models.}Appl. Math. Finance (2014).
%\bibitem{EA} {\sc Einstein. A.} {\it $\ddot{U}$ber die von der molekularkinetischen Theorie der W$\ddot{a}$rme geforderte Bewegung von in ruhenden Fl$\ddot{u}$ssigkeiten
%suspendierten Teilchen.} Annalen der Physik 17 (8), 549-560  (1905).
\bibitem{FoS}{\sc Forman. J. L and S\o{}rensen. M.} {\it The Pearson Diffusions: A Class of Statistically Tractable Diffusion Processes.} Scandi. J. Stat, Vol. 35: 438–465, 2008.
\bibitem{PJA} {\sc  Friz. P. K, Gatheral. J, Gulisashvili. A, Jacquier. A and Teichmann. J.}{\it Large Deviations and Asymptotic Methods in Finance.}  Springer Proceedings in Mathematics and Statistics, V. 110 (2015).
\bibitem{FOT}{\sc Fukushima. M, Oshima. Y and Takeda. M. } {\it Dirichlet Forms and Symmetric Markov Processes.} De Gruyter St. Math. 19 (2010).
\bibitem{GaSc}{\sc Garibaldi. U and Scalas. E. } {\it Finitary Probabilistic Methods in
Econophysics,} Cambridge University Press (Cambridge, UK, 2010).
\bibitem{GHLOW}{\sc Gatheral. J. Hsu. E. P. Laurence. P. Ouyang. C and Wang. T-H. } {\it Asymptotics of Implied Volatility in Local Volatility Models.} Math. Fin. 22, 591-620 (2012). 
%\bibitem{GVM}{\sc Gawarecki. L and Mandrekar. V.  }{\it Stochastic Differential Equations in Infinite Dimensions.}Springer Verlag Berlin Heidelberg, 2011.
\bibitem{GIS} {\sc Gihman. I.I and  Skorohod.A.V.} {\it Stochastic differential equations.} Springer-Verlag Berlin-Heidelberg. New York. 1972.
\bibitem{GST} {\sc Gottschalk. H , Smii. B and  Thaler. H.} {\it
The Feynman graph representation of general convolution semigroups
and its applications to L\'evy statistics}. Journal of Bernoulli
Society,14(2), 322-351 (2008).
\bibitem{GS}{\sc Gottschalk. H  and Smii. B. }{\it How to determine the law of the solution to a SPDE driven by a L\'evy space-time noise,} {Journal of Mathematical Physics}
 v.43. 1-22 (2007).
 \bibitem{GOS} {\sc Gottschalk. H, Smii. B and Ouerdiane. H.}
{\it Convolution calculus on white noise spaces and Feynman graph
representation of generalized renormalization flows.}
101-111pp.icmarp (2005).
\bibitem{GULI}{\sc Gulisashvili. A. } {\it Analytically Tractable Stochastic Stock Price Models. }
Lect. Notes. Maths, Springer (2012).
\bibitem{TH}{\sc Harg{\' e}. T.}{\it Borel summation of the small time expansion of the heat kernel. The scalar potential case.} \,\,arxiv:1301.7742v1. (2013).
\bibitem{TH1}{\sc Harg{\' e}. T.}{\it Borel summation of the semi-classical expansion of the partition function associated to the Schr\"{o}dinger equation.} (2013), https://hal.archives-ouvertes.fr/hal-00823084/file/ha8.pdf  .
\bibitem{TH2}{\sc Harg{\' e}. T.}{\it  Some remarks on the complex heat kernel on $\C^{\nu}$ in the scalar potential case,} arxiv: 1302.3512v1 [math-ph]. (2013)
\bibitem{TH3}{\sc Harg{\' e}. T.}{\it Borel summation of the heat kernel with a vector potential,} preprint (2013).
\bibitem{TH4}{\sc Harg{\' e}.T.}{\it A\,deformation formula for the heat kernel,}  (2013). https://pdfs.semanticscholar.org/cd43/d7b34a353296541cf1d53d267cc61428d087.pdf.
%\bibitem{HMN}{\sc Hairer, Martin, and Natesh S. Pillai. }{\it "Ergodicity of hypoelliptic SDEs driven by fractional Brownian motion." }Annales de l'Institut Henri Poincar{\'e}, Probabilit{\'e}s et Statistiques. Vol. 47. No. 2. Institut Henri Poincar{\'e}, 2011.
\bibitem{HkPS}{\sc Hida. T, Kuo. H-H, Potthoff. J and Streit. W.} {\it White Noise
An Infinite Dimensional Calculus.} Springer Netherlands. (1993).
\bibitem{JT}{\sc Jordan. R and Tier. Ch.  }{\it Asymptotic approximation to deterministic and stochastic volatilty models.} SIAM, J. Fin. Math. 2, 935-964 (2011).
\bibitem{Kan}{\sc Kanai. K. }{\it Short time asymptotics for fundamental solutions of partial differential equations.} Comm. Part. Diff. Eq. 2, 781-830 (1977).
\bibitem{KaShre}{\sc Karatzas, Ioannis, and Steven E. Shreve. }{\it Methods of mathematical finance.}Vol. 39. Springer Verlag, 1998.
\bibitem{Ki}{\sc Kifer. Y.I. }{\it Transition density of diffusion processes with small diffusion.} Theor. Prob. Appl. 21, 513-522 (1976).
\bibitem{Kol}{\sc Kolokoltsov. V.} {\it Semiclassical Analysis for Diffusions and Stochastic Processes.} Lect. Notes. Math. Springer (2000).
\bibitem{Kuso}{\sc Kusuoka. S and Osajima. Y.} {\it A remark on the asymptotic expansion of density function of Wiener functionals.}  JFA 255, 2545-2562 (2008).
\bibitem{La}{\sc Lamperti. J.\,\,\,}{\it A simple construction of certain diffusion processes.} J. Math. Kyoto Univ. 4, 161-170 (1964).
%\bibitem{LuPa}{\sc  Luschgy. H and Pag\`{e}s. G. }{\it Functional quantization of a class of Brownian diffusions: A constructive approach.} Stoch. Proc. Appl. 116, 310-336 (2006).
%\bibitem{LJ}{\sc ] Luo, Jiaowan. }{\it "Fixed points and exponential stability for stochastic %Volterra-Levin equations."}Journal of computational and applied mathematics 234.3 (2010): 934-940.
\bibitem{Lu}{\sc L\"{u}tkebohmert, Eva. }{\it An asymptotic expansion for a Black-Scholes type model. }Bull. Sci. Math. 128, 661-685 (2004).
%\bibitem{LOK} {\sc L${\o}$kka. A, ${\O}$ksendal. B and Proske. F.}
%{\it Stochastic partial differential equations driven by L\'evy
%space-time white noise.} Ann. Appl. Prob. 14, 1506-1528 (2004).
\bibitem{Ma}{\sc Mandelbrot, Benoit B. }{\it Fractals and scaling in finance: Discontinuity and concentration.}Springer Verlag, 1997.
%\bibitem{MBPR} {\sc Meyer-Brandis. T and Proske. F} {\it Explicit
%representation of strong solutions of SDEs driven by infinite
%dimensional L\'evy processes.} J. Theor. Prob. 23, 301-314 (2010).
\bibitem{MM} {\sc Milan. P and Miroslav. k.} {\it Stochastic equations for simple discrete models of epitaxial
growth.} Phys. Rev. E.  V.54 (4), (1996).
\bibitem{Mo}{\sc  Molchanov. S.A. }{\it Diffusion processes and Riemannian Geometry.}  Math. Surv. 30, 1-63 (1975).
\bibitem{MMH}{\sc  M\o{}ller. J.K and Madsen. H.}{\it From State Dependent Diffusion to Constant Diffusion in Stochastic Differential Equations by the Lamperti Transform.} Tech. Univ. of Denmark, DTU Informatics, 25 p. 2010.
%\bibitem{MG}{\sc  Mugnolo. D}{\it Gaussian estimates for a heat equation on a network, Netw. Heter.}Media 2, 55-79, 2007.
%\bibitem{PSJ}{\sc Peszat, Szymon, and Jerzy Zabczyk. }{\it  Stochastic partial differential equations with L{\' e}vy noise: An evolution equation
%approach} Vol. 113. Cambridge University Press, 2007.
\bibitem{SA}{\sc  Sengupta, Ambar N. }{\it "Pricing derivatives." }Ed. Mcgraw-Hill Professi (2005).
%\bibitem{DN} {\sc Nualart. D.} {\it Analysis on Wiener space and anticipating stochastic calculus, Lecture notes in Mathematics, V 1690/1998, 123-220, 1998. DOI: 10.1007/BFb0092538.}
\bibitem{OKS} {\sc $\O${ksendal}. B} {\it Stochastic Differential Equations: An introduction with Applications.} Springer -Verlag Berlin Heidelberg, 6th Edi. (2010).
%\bibitem{OU} {\sc Ornstein. L. S and Uhlenbeck G. E.} {\it On the Theory of the Brownian motion} Phys. Rev. 36, 823-841 (1930).
\bibitem{RS} {\sc Reed. M and Simon. B. }{\it IV Analysis of Operators.} Elsevier Science Publishing Co Inc. V4. 1978.
\bibitem{Rez} {\sc Rezende. J.} {\it  The method of stationary phase for oscillatory integrals on Hilbert spaces.} Comm. Math. Phy. V 101, 2, pp. 187-206 (1985).
%\bibitem{Ru} {\sc Ruelle. D. }{\it Statistical mechanics -- rigorous results.} Benjamin, London / Amsterdam / Don Mills (Ontario) /
%Sydney / Tokyo 1969.
%\bibitem{KS} {\sc  Sato. K.I.} {\it L\'evy processes and infinite divisibility.} Cambridge University Press, 1999.
%\bibitem{Se} {\sc Sengupta. A.} {\it Pricing derivatives." Ed. Mc-Graw-Hill Professi (2005).}
% \bibitem{ST} {\sc Shiga. T.}  {\it A recurrence criterion for Markov processes of Ornstein-Uhlenbeck type}, Prob. Th. Rel.Fields. 85, pp. 425-447 (1990).
 %\bibitem{BS} {\sc Smii. B.}  {\it A large diffusion expansion for the transition function of L\'evy Ornstein-Uhlenbeck processes.} Appl. Math. Inf. Sci. 10, No.4, pp.1-8(2016).
 %\bibitem{BS1} {\sc Smii. B.}  {\it A Linked Cluster Theorem of the solution of the generalized Burger equation}, Applied Mathematical Sciences.(Ruse),
% vol. 6, no. 1, pp. 21-38, (2012).
%\bibitem{SM} {\sc Smoluchowski. M.}  {\it Zur kinetischen Theorie der Brownschen Molekularbewegung und der
%Suspensionen.} Annalen der Physik 21 (14), 756-780 (1906).
 \bibitem{SO} {\sc Sokal. A. D.} {\it An improvement of Watson's theorem on Borel summability}, J. Math. Phys. 21.2, 261--263 (1980).
 \bibitem{TA} {\sc Takahashi. A.\,} {\it Asymptotic Expansion Approach in Finance.} Large Deviations and Asymptotic Methods in Finance, Springer Proceeding in Mathematics and Statistics. V. 110. DOI 10.1007/978-3-319-11605-1 ,  (2015).
%\bibitem{TL} {\sc Taufer. E and Nikolai. L.} {\it Simulation of L\'evy-driven Ornstein-Uhlenbeck processes with given marginal
%distribution,} Computational Statistics and Data Analysis, Elsevier,
%vol. 53(6), pages 2427-2437,(2009).
\bibitem{Ta}{\sc Tuckwell, Henry Clavering. }{\it Stochastic processes in the neurosciences. }Vol. 56. Society for industrial and applied mathematics, 1989.
\bibitem{Tub}{\sc Tuckwell, Henry C. }{\it "Analytical and simulation results for the stochastic spatial Fitz Hugh-Nagumo model neuron." }Neural computation 20.12 (2008): 3003-3033.
\bibitem{UMN}{\sc Uchida, Masayuki, and Nakahiro Yoshida. }{\it "Information criteria for small diffusions via the theory of Malliavin-Watanabe." }Statistical Inference for Stochastic Processes 7.1 (2004): 35-67.
\bibitem{UM}{\sc Uemura. H.}{\it On a short time expansions of the fundamental solution of heat equations by the method of Wiener functional.} J. Math. Kyoto Univ., 27, 417-431 (1987).
\bibitem{Vas}{\sc Vasicek. O.} {\it An equilibrium characterization of the term structure.} J. Financ. Ec. 5, 177-188 (1977). 
%\bibitem{WA}{\sc Walsh, John B.  }{\it An introduction to stochastic partial differential equations.  }Springer Berlin Heidelberg, 1986.
\bibitem{Wafa}{\sc Wang. T-H and Gatheral. J.}{\it Implied Volatility from Local Volatility: A Path Integral Approach.} Springer Proceedings in Mathematics and Statistics (2015).
\end{thebibliography}

\documentclass[12pt,twoside]{article}

\usepackage{setspace}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{graphicx,color}
\usepackage{pstricks}
\usepackage{pst-text}
\usepackage{pst-tree}
\usepackage{multido}
\def\R{{\mathbb R}}
\def\N{{\mathbb N}}
\def\C{{\mathbb C}}

\newtheorem{thm}{Theora}[section]
\newtheorem{theo}[thm]{Theorem}
\newtheorem{hypothesis}[thm]{Hypothesis}

\newtheorem{rem}[thm]{Remark}

\newcommand\dint{\displaystyle\int}
\newcommand\ds{\displaystyle\sum}

\newcommand\dlim{\displaystyle\lim}

\psset{xunit=1cm,yunit=1cm,runit=1cm,shortput=tab}
\begin{document}
\pagestyle{myheadings} \markboth{ Albeverio.S and Smii.B }{Borel summation of the heat kernel} \thispagestyle{empty}

\title{Borel summation of the small time expansion of some SDE's driven by Gaussian
white noise}
\author{Sergio Albeverio\footnote{Dept. Appl. Mathematics, University of Bonn, HCM, BiBoS, IZKS; KFUPM Dhahran;  CERFIM, Locarno. {\tt albeverio@uni-bonn.de}}\,\,\, %
and\,\,
  Boubaker Smii\footnote{ King Fahd University of
Petroleum and Minerals, Dept. Math. and Stat., Dhahran 31261, Saudi
Arabia. \hspace*{5mm}{\tt boubaker@kfupm.edu.sa}}}

\maketitle
 \begin{abstract} 
We consider stochastic differential equations with a drift term of gradient type and driven
by  Gaussian white noise on $\R^d$. 
Particular attention is given to the kernel  $p_t,\,t\geq 0$ of the transition semigroup associated with the solution process.\\
 Under some rather strong regularity and growth assumptions on the coefficients, we adapt previous work by Thierry Harg\'e on Schr\"{o}dinger operators and prove that the small time asymptotic expansion of $p_t,\,t> 0$ is Borel summable.\\
 We also briefly indicate some extensions and applications.

\end{abstract}
\noindent {\bf Key words:} {\em SDEs, asymptotic expansions,  Borel summability.}\vskip0.2cm

\section{Introduction and motivation}
In the description of the dynamics of many phenomena studied in natural sciences, engineerings as well as in economics and social sciences the use of stochastic differential equations has been very extensive, in order to cope both with deterministic and random effects. Very often solutions of these equations are obtained by complicated constructions and it is difficult to extract more concrete information about related quantities of interest. Then one has recourse to numerical or computational methods, that involve suitable approximations, or,often alternatively, one develops asymptotic methods, like expansions with respect to parameters or variables appearing in the equations.\\
In the present paper we shall look at one particular type of expansion, namely small time expansions for stochastic differential equations for diffusion process, with Brownian motion as driving process.\\
The usefulness of such expansions has been pointed out very early in many contexts, including differential geometry, analysis, quantum mechanics and statistical mechanics, see, e.g. \cite{ACDP, AlDiPMa, ALDIBOUB, ALDIBOUB1, AGoWu, AHk, ALEBOU, ALRS, Az, Ba, BE, BBF, CL, Co1, Co2, GHLOW, GULI, HkPS,  JT, Kan, Ki, Kol,  Kuso, Lu, Mo, TA, Ta, Tub, UMN, UM} and references therein.\\ 
Applications to various areas including physics, neurobiology, see, e.g. \cite{ABrMa, AlGYo, AKKR, GST, GS, GOS, MM, Ta} and mathematical finance, see,e.g. \cite{GULI} have been also provided. For the latter area let us mention particularly that besides well known models discussed in various textbooks, also models with  inclusion of interacting assets with different types of noise have also been studied in \cite{ALiMa, ALBO, ASte, Lu}.\\
Advanced statistical methods have also been applied to finance in different directions, from econometrics to methods inspired by statistical mechanics, see, e.g., \cite{Ma}, \cite{GaSc}.\\
From a mathematical point of view we shall
develop here first results concerning the study of important properties of
solutions for a class of  SDEs driven by an additive Gaussian white
noise. At the end of the paper we mention a couple of applications.\\
Here we shall concentrate on the class of SDEs of the form:
\begin{equation}\label{e12}
\left\{ \begin{array}{ll} dX_{t}=\beta (X_{t})\,dt+\sigma (X_{t})\,
d\,W_{t}&,\,\, t\geq 0\\ X_0= x^0,\,\,\,x^0\,\in \, \R^d,
\end{array}
\right.
\end{equation}
where $X_t=(X_t)_{t\geq 0}$ is a stochastic process taking values in
a finite dimensional space $\R^d$,  while $\beta$ is a vector field from\,$\R^d$ to $\R^d$ and $\sigma$ is a $d\,\times\,d$-matrix with real coefficients depending on the space variable in $\R^d$, $W_t$ is a standard Brownian motion (Wiener process in $\R^d$).\\
Under suitable regularity and growth assumptions on $\beta$ and $\sigma$, existence and
uniqueness of solutions are known, see, e.g. \cite{GIS, OKS}. \\
Moreover, under additional regularity assumptions the transition semigroup of the solution process $X_t, t\geq 0,$ has a density (with respect to Lebesgue measure on $\R^d),\, p_t(x,y)\,,t\geq 0, x,y\,\in\,\R^d.$ For short we recall  that the transition kernel $p_t$ satisfies then the following Kolmogorov forward equation (Fokker-Planck equation)(the space derivatives being understood with respect to $x$).
\begin{equation}\label{e13}
\left\{ \begin{array}{ll} {d \over \partial t}p_{t}(x,y)=\,\Delta[a(x) \,p_t(x,y)]-\nabla[\beta(x)\,p_t(x,y)]&,\,\, t > 0\\ \dlim_{x \longrightarrow 0}p_0(x,y)\,dx=\delta_0(x-y),
\end{array}
\right.
\end{equation}
 with $a:={1 \over 2}\,\sigma\,\sigma^t,\,\delta_0$ being Dirac measure at the origin, see, eg., \cite{GIS}, \cite{OKS}.\\
 In this paper we concentrate on the case where $\sigma$ is proportional to the unit $d\times d$-matrix\,${\bf 1}$. Only in the last section we consider the case with general $\sigma$ but $d=1$, reducing it by a suitable time change transformation to the case $\sigma=1$. \\
 We discuss the asymptotic expansion of $p_t$ for $t\downarrow\,0$ and its Borel summation.
We  rely essentially on a setting and results contained in important work done by Thierry Harg\'e \cite{TH, TH1, TH2, TH3, TH4} for the study of the Schr\"{o}dinger equation for analytic potentials and its corresponding diffusion equation. For this we need to assume throughout our paper that $\beta$ is a gradient field.\\
Let us now describe the single sections of this paper:\\
In the section 2 we will introduce the basic technical assumptions, provide definitions of Borel summability and present the important results obtained by T. Harg\'e.\\
Section 3 will be reserved to the study of Borel summation of the small time asymptotic expansion of the above transition kernel $p_t$.\\
Section 4 is devoted to some comments on extension to the case $\sigma\neq 1 $, at least for $d=1$, and on some applications.\\
\section{Assumptions, definitions and the results by T. Harg\'e}
In this section we will outline the assumptions needed for adopting to our setting results obtained by T. Harg\'{e}, see. \cite{TH, TH1, TH2, TH3, TH4} for the time dependent Schr\"{o}dinger equation.
\begin{hypothesis}\label{hp:A+F}
Let us consider "potential functions" $V$ from $\R^d$ into $\R$ of the form $V(x):=-\dint e^{i\,x\,\xi}\, \mu_V(d\xi),\,x\,\in\,\R^d$, for some complex Borel measure $\mu_V$ on $\R^d$ such that $\mu_V(-A)= \mu_V(A)$ for any Borel subset $A$ of $\R^d$ (this implies indeed that $V$ is real valued: this property will be indicated by saying that $\mu_V$ is symmetric).\\
(Here and in the whole paper $\R^d$ is Euclidean $d-$dimensional space for any natural number $d$.)  
        \begin{enumerate}
        \item[]
             \item $V$ is assumed moreover to be such that $-\Delta +V \geq E_0$ for some constant $E_0\geq 0$ and there exists a strictly positive $C^1(\R^d)$ function $ \,\phi\,\in\, L^2(\R):=L^2(\R, dx),$ satisfying $(-\Delta +V)\phi=E_0\,\phi$ for some $E_0\geqslant 0.$ We can always assume that $E_0=0,$ otherwise we would change $V$ in $\tilde{V}=V-E_0$ (with corresponding $\mu_{\tilde{V}}(dx)=(\mu_V-E_0\,\delta_0)(dx)$, with $\delta_0(dx)$ the Dirac's measure at zero). In this case, i.e., when  $E_0=0,\, -\Delta +V$ is  non negative on $C^2_0(\R^d)$ in $L^2(\R^{d})$. \\
           
         \item Let $\beta:\R^d\,\rightarrow\,\R^d$ be  such that $\beta(x):=2\,\phi^{-1}(x)\,\nabla\,\phi,$  then the unitary map $U_{\phi}$ from $L^2(\R^{d}, \phi^2\,dx)$ to $ L^2(\R^d) $ given by multiplication by $1 \over {\phi}$ transforms the operator $H=-\Delta-\beta.\nabla$ acting (on the dense domain) $\phi^{-1} C^2_0(\R^d) $ in $L^2(\R^{d}, \phi^2\,dx) $ into the operator  $-\Delta\,+V$ acting (on the dense domain) $ C^2_0(\R^d) $ in $\,L^2(\R^{d},\,dx):=L^2(\R^{d})$, in the sense that
                
         \begin{equation}
         H\,\phi^{-1}\,f=\,\phi^{-1}(-\Delta+V)\,f\,\,\,\,\,\,\forall\,f\,\in\,C^2_0(\R^d).
         \end{equation}
   This transformation is called ground state transformation, see, e.g., \cite{AHKS, Al, AlMa, RS}      
      \item Under some regularity assumptions on $V,$ and consequently on  $\phi,\,$ $-\Delta\,+V\,$ is essentially self adjoint and positive on $C_0^2(\R^d)$ in $L^2(\R^{d})$, and correspondingly for $H$ on $ \phi^{-1} C^2_0(\R^d)$ in $L^2(\R^{d}, \phi^2\,dx)$. Hence their closures, called again $H$ resp. $-\Delta+V $, are generators of strongly continuous contraction semigroups $e^{-tH}$ resp. $e^{-t(-\Delta+V)}$ in $L^2(\R^{d}, \phi^2\,dx)$ resp. $L^2(\R^{d})$  and 
          \begin{equation}
                   e^{-tH}\,\phi^{-1}\,f=\,\phi^{-1}(e^{-t(-\Delta+V)})f\,\,\,\,\,\,\forall\,f\,\in\,C^2_0(\R^d), t \geq 0.
          \end{equation}
  \end{enumerate}           
 \end{hypothesis}
  The relation between $V$ and $\phi$ can be expressed through $\beta$ by the formula
  \begin{equation}
  V={1 \over 2}\,({\mid\beta\mid^2 \over 2}+\,div \beta)\,\,.
  \end{equation}
 In the following concepts like asymptotics expansions, Borel summability, Borel transform and Borel sum will be needed, we recall them briefly following \cite{TH} (see also the related papers \cite{TH1}, \cite{TH2}.)\\
 For any  $\kappa,\, T> 0$ we define ${\cal P}_{\kappa,\,T}$ as the set of complex-valued functions $f$ of a complex variable $t$ which are analytic on ${\tilde D}_{T}:=\{ z\,\in\,\C\,\mid Re{1 \over z} > {1 \over T}\}$ (the open disk of center ${T \over z}$ and radius ${T \over z}$ ) and such that there exist sequences $(a_k,\,R_{k}),\,k\,\in\,\N_0,$ of complex analytic functions on ${\tilde D}_{T}$ such that:
 \begin{enumerate}
 \item For every $r \in\,\N, k, T>0$ and $t\,\in\,{\tilde D}_{T} $ it holds that
 \begin{equation}
  f(t)=a_0+a_1\,t+\cdot\cdot\cdot+a_{r-1}\,t^{r-1}+R_r(t).
 \end{equation}
 We then say that $f \,\in\,{\cal P}_{\kappa,\,T} $ with coefficients $\{a_k\}$ and remainders $R_k,\, k\,\in\,\N_0$.\\
 
 \item For every $\bar{\kappa} < \kappa,\,\, \bar{T}< T $ there exists $K>0$ (depending on $f, r$) such that for every $r\,\in\,\N_0$ and $t\,\in\,{\tilde D}_{T}$ the following estimates hold
 \begin{equation}
 \mid R_r(t) \mid \leq K\, {{r!} \over {\bar{\kappa}}^r } \,\mid t\mid^r.
 \end{equation}
 
 \item For any  $\kappa,\, T> 0$ we define ${\cal Q}_{\kappa,\,T}$ to be the set of functions $g$ such that $g$ is complex analytic on ${\tilde S}_{\kappa}:=\{ z\,\in\,\C\,\mid d(z,\,[0,+\infty))< \kappa   \}$ and for every $\bar{\kappa} < \kappa,\,\, \bar{T}< T $ there exists $\tilde{\kappa} $ such that for every $\tau\,\in\,{\tilde S}_{\bar{\kappa}}$ the following estimate holds
  \begin{equation}\label{KAS}
  \mid g(\tau) \mid \leq K_g\, \exp({\mid \tau \mid}/{ \bar{T}}),
  \end{equation}
 for some constant $K_g$ (depending on $g$), with $d(z,A):=inf\{\mid z-w \mid,\,w\,\in\,A \},\,\,A$ a subset of $\C$.\\
The following theorem goes back to Nevanlinna, who in turn improved a theorem of Watson, see \cite{SO}, \cite{Rez} .

\begin{theo}\label{ABDEL}
If\, $f \in {\cal P}_{\kappa,\,T}$, for some $\kappa,\, T> 0$, with coefficients $\{a_k\}$ and remainders $\{ R_k\}\,,\,k\,\in\,\N_0$, then 
 \begin{equation}\label{NA}
 B(f)(\tau):=\ds_{r=0}^\infty {{a_r} \over {r!}}\,\tau^r
 \end{equation}
converges absolutely for  $ \tau \in \R$ and admits an analytic continuation $g$ to ${\tilde S}_{\kappa}$ which belongs to ${\cal Q}_{\kappa,\,T}$. Moreover, if $h$ is a function in ${\cal Q}_{\kappa,\,T}$ then 
\begin{equation}\label{NAW}
 f_h(t):={1 \over t}\dint_0^\infty h(\tau)\,e^{-{\tau \over t}}\,{d\tau}
 \end{equation}
is in ${\cal P}_{\kappa,\,T},\,$ for any $t \in\, {\tilde{D}}_{T}$. In particular, if $h=g$ then $f_g$ coincides with $f$.
\end{theo}
\begin{rem}
\begin{enumerate}
\item  $f_g$ is the Laplace transform of $g$ and $B(f)$ as defined in (\ref{NA}) is the Borel transform of $f$.\\
\item If $f \in\,{\cal P}_{\kappa,\,T} $, then the coefficients $a_r$ are uniquely determined and their knowledge suffices to recover $f$ through the analytic continuation $g$ of its Borel transform  $B(f)$ as given by (\ref{NA}). (Thus $f$ is then the Laplace transform of its Borel transform $B(f)$, i.e., $g$ satisfies $Q_{k, T}$ and for $t \in {\tilde{D}}_{T}:$
\begin{equation}
f(t)=f_g(t),
\end{equation}
for $g=B(f)$) (and $B(f)$ only depends on the $a_r\,r\,\in\,\N_0$, by (\ref{NA})).

\end{enumerate}
\end{rem} 
Finally let us define Borel summability and Borel sum.\\
Let $\tilde{a}_r,\,r\,\in\,\N_0$ be complex numbers. A formal power series $\ds_r \tilde{a}_r\,t^r,\,t\,\in\,\C$ is said to be Borel summable if there exist $\kappa,\, T> 0$ and a function $f \in {\cal P}_{\kappa,\,T}$ with  coefficients $a_r,\,r\,\in\,\N_0$ s.t. for every
 $r\,\in\,\N_0,\,a_r=\tilde{a}_r.\, $ 
 $f$ is then called Borel sum of the formal power series $\ds_r \tilde{a}_r\,t^r$ in ${\tilde{D}}_{T}$.

\end{enumerate}  
                   
We have from \cite{TH} (Theorem 3.1)(restricting ourselves to the case of a complex-valued $\mu_V$ (instead of the matrix-valued as assumed in \cite{TH}, since it suffices for the applications that we have in mind, see. Section 4,) the following result:  
\begin{theo}(Harg\'e, \cite{TH1}) \label{PA}
 Let $V$ be  as in Hypothesis 2.1, and assume in addition that $\dint e^{a\,\mid\xi\mid^2}\mid\mu_V
\mid(d\xi)< \infty , \, $  for some $a > 0$ (this is a regularity assumption on $V$).\\
Let $p_t(x,y)\,,t \geq 0,\,x,y\,\in\,\R^d$ be the kernel of the strongly continuous contraction semigroup $e^{t(\Delta-V)}$ in $L^2(\R^{d})$.
\begin{enumerate}
\item 
Let , for $t > 0, \,\tilde{p}_t(x,y):=(4\,\pi\,t)^{d \over 2}\,e^{{(x-y)^2 \over {4\,t}}}\,p_t(x,y) $. Then for every $x,y\,\in\,\R^d, $ $\tilde{p}$, as a function of $t$, has a Borel transform $B(\tilde{p})(\tau, x, y)$ defined for $\tau \,\in\,\R$ by (\ref{NA}) (with $a_r$ the coefficients of the expansion of $\tilde{p}_t$ in powers of $t$), which is analytic in $(\tau,\,x,\,y)$ on $\mathbb{C}^{1+2d}$.\\
\item
Let $S_{\kappa}:=\{ z\,\in\, \mathbb{C}\,| \,\,\,\, | Im\,z^{1 \over 2}|^{2}\,\,< \kappa \}$, (where for any $z\,\in\,\C$ of the form $z=\mid z \mid\,e^{i\,\theta},\, \theta \,\in\, (-\pi,\,\pi]$ we denote $z^{1 \over 2}=\mid z \mid^{1 \over 2}\,e^{i\,{\theta \over 2}} $)  , for a given $\kappa > 0$ (this is a parabola which contains $\tilde{S}_{\kappa}$, as defined just before (\ref{KAS}) ). Then  for every $(\tau, x, y)\, \in\, S_{\kappa}\,\times\, {\mathbb{C}_R^{2d}}$, with ${\mathbb{C}_R^{2d}}:=\{ (x ,y)\,\in\,\mathbb{C}^{2d},\,  | Im\,x| < R,\,\, | Im\,y|< R\},$ for some given constant $ R> 0$, we have $|B(\tilde{p})(\tau,\,x,\,y)|\leq \exp(C\,|\tau|^{1\over 2}),$ with 
\begin{equation}C := 2\,e^{{\kappa \over a}}\Big(\dint_{\R^d}\exp({a \over 2}\xi^2+ R\mid \xi \mid) \mid \mu \mid(d \xi) \Big)^{1 \over 2},
\end{equation}
 (using the same symbol $B(\tilde{p})$ for the analytic continuation of $B(\tilde{p})$ from $\R \times \C^{2d}_R$ to $S_{\kappa} \times \C^{2d}_R$ ).\\
Moreover,\, for any $(x ,y)\,\in\,\mathbb{C}_R^{2d}$,\, the small time expansion of\,\, $\tilde{p}_t(x,y)=(4\,\pi\,t)^{d \over 2}\,e^{{(x-y)^2 \over {4\,t}}}\,p_t(x,y), t > 0,$ in powers of $t$ exists, $\tilde{p}_t $ is in ${\cal P}_{\kappa,\,T},$ (with ${\cal P}_{\kappa,\,T}$ as defined before Theor. \ref{PA}), is Borel summable and its Borel sum is equal to $\tilde{p}_t$.
\end{enumerate}
\end{theo}
\begin{proof}
 The proof is provided in \cite{TH1}.
 \end{proof}
\begin{theo}(Harg\'e) \label{PAP}
Let $\omega \,\in\,\C,\,V$ as in Theorem \ref{PA}. Let $u(t, x, y),\,t >0,\, x,y\,\in\,\R^d$ be the solution of 

\begin{equation}
 \partial_t\,u(t, x, y)=(\Delta_x-{{\omega^2 \over 4}}\,x^2)\,u(t, x, y)-V(x)\,u(t, x, y),\,\,x\,\in\,\R^d,
\end{equation}

with $\dlim_{t \longrightarrow\,0} u(t, x, y) = \delta(x-y) $ in the distributional sense. $u(t, x, y), t > 0 $ is then, for $\omega \,\in\,\R,$ the density of the kernel of the operator $e^{t(\Delta-W)}$ in $L^2(\R^d)$ with $W(x)=V(x)+{{\omega^2 \over 4}}\,x^2$. Then  for any $r \in \N_0$, there exists $T> 0$ and analytic functions $a_r$ on $\mathbb{C}^{2d}$ and $R_r$ on $D_T^{+} \times \mathbb{C}^{2d},$   with\\ $D_T^{+}:= D_T \cap \C^{+} ,\,\,\,D_T:= \{ z\,\in\, \mathbb{C}\,| \,\,\,\, | z|\,< T \}$, such that 
\begin{equation}\label{IS}
u(t, x, y)= (4\,\pi\,t)^{-{d \over 2}}\,e^{-{(x-y)^2 \over {4\,t}}}\,\Big( 1+a_1(x,y)\,t+\cdot\cdot\cdot+a_{r-1}(x,y)\,t^{r-1}+R_r(t,x,y)\Big),
\end{equation}
for every $r\, \in \N, \,t\,\in\,D_T^{+},\, (x,y)\,\in\, \mathbb{C}_R^{2d}.$\\
Moreover, for each $R> 0$, there exist $ K,\,\kappa >0$ such that 
\begin{equation}\label{IS1}
 \mid R_r(t, x, y) \mid \leq K\, {{r!} \over {{\kappa}}^r } \,\mid t\mid^r,
 \end{equation}
for every $r\,\in\,\N_0,\,t\,\in\, D_T^{+},\,(x,y)\,\in\, \mathbb{C}_R^{2d}$.\\
The expansion for\, $u(t,x,y)\,(4\,\pi\,t)^{{d \over 2}}\,e^{{(x-y)^2 \over {4\,t}}}:=v(t, x, y)$ (obtained from the one in (\ref{IS}) multiplying both members by $(4\,\pi\,t)^{{d \over 2}}\,e^{{(x-y)^2 \over {4\,t}}} $), as a function of $t$, constitutes an asymptotic series in powers of $t,\, \sum_{j=0}^{r-1}\, a_j(x, y)\,t^j\,+ R_r(t, x, y), $ with $a_0\equiv 1,$ that has a Borel transform belonging to ${\cal Q}_{\kappa, T}$, for any $x, y \in \mathbb{C}_R^{2d}.$ The Laplace transform (\ref{NAW}) of this Borel transform coincides with $v(t,x,y)$. The asymptotic series in powers of $t$ constituting the above small time expansion of $v(t,x, y)$ is Borel summable to $v(t, x, y)$.
\end{theo}
\begin{proof}
 For the proof we refer to  \cite{TH}, Theor. 3.5.
 \end{proof}
\begin{rem}
\begin{enumerate}
\item If $\omega\,\in\,\R, u(t,x,y)$ is the kernel of an analytic semigroup, the one given by the density of the kernel of $e^{t(\Delta-W)}$ in $L^2(\R^d)$.
\item For $\omega \in i\R,$ \cite{TH} has also provided a uniqueness result for analytic semigroups related to $u$.
\end{enumerate}
\end{rem}
 \section{Borel summation of the transition kernel for the SDE with additive noise}
 Following Theorems \ref{PA} and \ref{PAP} and the assumptions on $V$ given in the previous section we have the following small time expansion of (a quantity simply related to) the kernels giving the transition semigroup for the solution process of the SDE (1) with $\sigma \equiv {\bf 1}$.
 
 \begin{theo} \label{PAAP}
 Consider the stochastic differential equation
 \[\displaystyle  \left\{\begin{array}{lll} dX_{t}=\beta\, (X_t)\,dt+dW_{t}\\ \displaystyle x_0=x^0,\,\,x^0 \in \mathbb{R}^d\\ \end{array}\right.\]
 
 where the drift $\beta$ is a gradient field such that
 $\beta(x)=2\psi^{-1}\,(x)\nabla \psi\,(x),$ for a primitive smooth
 function $\psi$ of the form \begin{equation} \label{TUN}
 \displaystyle \psi{(x)}=c_{\varphi}\,.\,\varphi(x)\,
 \frac{e^{-\omega\frac{| x |^2}{4}}}{(2\pi\,{2 \over {\omega}})^{d/2}},
 \end{equation}
 
 where $\varphi > 0$ is a smooth function for $\mathbb{R}^d$ to
 $\mathbb{R}_{+}, \,\,\omega^2>0,$ such that $\displaystyle
 {\mid\varphi\mid^2}\,e^{-{\omega\over 2}\mid\cdot\mid^2)}\in L^1\,(\mathbb{R}^d)$.\\ $c_{\varphi}$ is the normalization constant given by
 \begin{equation} 
 \displaystyle c_{\varphi}= \left(\int_{\mathbb{R}^d} \varphi^2(x) \frac{e^{-{\omega\over 2}|x|^2}}{\left(2\pi\frac{2}{\omega}\right)^{d}}\,\,dx\right)^{-1},
 \end{equation}
 so that $\displaystyle \int_{\mathbb{R}^d}|\psi^2|(x)\,dx=1.$\\ 
   Set $\beta_{\varphi}(x)=2\,\varphi^{-1}(x)\nabla \varphi (x)$ and
 assume $V_{\varphi}(x):={1 \over 4}\mid\beta_{\varphi}(x)\mid^2+ {1 \over 2}\, div\,\beta_{\varphi}(x)$
 satisfies the assumption $2.1$, and
 \begin{equation} 
 V_{\varphi}(x)=-\int_{\mathbb{R}^d}\,e^{i xy}\, \, \, \mu_{V}\, \, (dy), \end{equation}
 for some bounded symmetric complex measures $\mu_{V}$ on
 $\mathbb{R}^d,$ such that (as theorem \ref{PA}, for $V=V_{\varphi})$,\,\,\,
 $\displaystyle \int_{\mathbb{R}^d}e^{a|y|^2}\,
 |\mu_{V}|\,(dy)<\infty,$ for some $a>0$.\\
  Then  the following properties hold:\\
 
 (i) The unique solution process $X_t$ of (1) is a diffusion process
 with drift $\beta(x)=\beta_{\varphi}(x)-\omega\,x,$ and invariant
 symmetrizing probability measure $\nu(dx)=\psi^2\,(x)\,dx.$\\
  The corresponding Markov transition semigroup $p_t, t>0$ has a kernel with
 density $k(t,x,y),\,t>0,\, x,y\in\mathbb{R}^d$, such that
 $(p_t\,f)\,(x)=\int_{\mathbb{R}^d}\, k(t,x,y)\, f(y)\,dy,$ for any
 bounded $C_0\,(\mathbb{R}^d)$- function $f$. $p_t$ is the symmetric
 Markov $C_0$ semi-group in $L^2(\mathbb{R}^d,\, \nu)$ associated
 with the classical Dirichlet form $\varepsilon_{\nu}$ given by $\nu$,
 i.e.
 \begin{equation}
 \varepsilon_{\nu}\,(f,g)=\int_{\mathbb{R}^d}\, \nabla f
 \cdot \nabla g\, d\nu.
 \end{equation} 
 The generator $L_{\nu}$ of $p_t$ is the self-adjoint operator in
 $L^2(\mathbb{R}^d, \nu)$ given on smooth function $f$ of compact
 support by $L_{\nu}\,f(x)=(\triangle+\beta_{\varphi}(x)\cdot \, \nabla\,-
 \omega\, x \nabla)\,f(x).$\\ The operator $-L_{\nu}$ is unitary equivalent with the
 self-adjoint operator
 $H=-\triangle+V_{\varphi}(x)-{\omega \over 2}\,x\,\beta_{\varphi}(x)+\frac{\omega^2}{4}|x|^2-{d \over 2}\,\omega$ acting in $L^2\, (\mathbb{R}^d).$ 
 
 (ii) One has $k(t,x,y)=\psi^{-1}(x) \, u (t,x,y)\, \psi(y),$ with
 $u(t,x,y)$ as in Theorem \ref{PAP} (with $\omega >0$). $u(t, x, y)$ satisfies the
 asymptotic expansion given by (\ref{IS}) and (\ref{IS1}), and consequently $k(t, x, y)$ has a corresponding asymptotic expansion, given by (\ref{IS}), (\ref{IS1}) multiplied by $\psi^{-1}(x) \,\psi(y) $.\\
  The quantity
 \begin{equation}\label{HI}
  \tilde{k}(t,x,y):=(4\,\pi\,t)^{d/2}\, e^{{(x-y)^2} \over {4\,t}}\, k(t,x,y), 
 \end{equation}
 
 which we call modified kernel function for the solution process
 $X_t$ to (1), has, as a function of $t$, an asymptotic expansion in
 powers of t, its Borel transform belonging to $Q_{\kappa,T},$ for any $x,y
 \in\mathbb{C}^{2d}_{R}, \, R>0$( with $Q_{\kappa,T}$ as in point 3 of Hypothesis 2.1 ). The asymptotic expansion for $\tilde{k}(t,x,y)$ is given by\,\,\, $\psi(x)\Big[ 1+ \ds_{j=1}^{r-1} a_j\,\psi(x)(1+ \ds_{j=1}^{r-1} a_j(x ,y)\,t^j+\,R_r(t,x,y))\Big]\,\psi^{-1}(y),$ with the $a_j,\,j=1,...,r-1\,$  as in (\ref{IS}).\\
 The  Laplace transform of this Borel transform (i.e. of $k(t, x, y)$) coincides with the
 function given in (\ref{HI}), which constitutes then the Borel sum of the asymptotic
 expansion in powers of $t$ for the modified transition kernel $\tilde{k}(t,x,y)$ for
 the solution of the SDE.
  \end{theo}
 \begin{proof}
 i) The proof uses essentially the theory of Dirichlet forms and the
 regularity assumption on $\psi$, see, e.g. \cite{Al}, \cite{AHKS}, \cite{FOT}.\\
  Let us look at the Dirichlet form $\varepsilon_{\nu}$ defined on
 $C^{\infty}_{0}\, (\mathbb{R}^d)$. By integration by parts we have
 for $f,g \in C^{\infty}_{0}\, (\mathbb{R}^d):$
 \begin{equation}
 \varepsilon_{\nu}\, (f,g)=\int \nabla f \cdot \nabla g\, d\nu=-\int f
 \triangle \,g \, d\nu-\int f \, \beta_{\psi}\nabla g \, d\nu
 ,\end{equation}
  where \begin{eqnarray}\label{ZIT} \beta_{\psi}(x)&=& 2 \, \nabla \ln\, \psi(x)\nonumber\\&=& 2\nabla \ln\,
 \varphi(x)+2\nabla \ln c_{\varphi} \frac{e^{-\omega{|x|^2 \over 4}}}{(2\pi
 \frac{2}{\omega})^{d\over 2}}\nonumber\\&=& \beta_{\varphi}(x)-\omega\,x,
 \end{eqnarray}
 where we used the definition of $\psi, \beta_{\psi}$ and $\beta_{\varphi}$. Hence we found the given expression for $L_{\nu}.$ The function identically equal to ${\bf 1}$ is the domain of the positive operator $L_{\nu}$ and we have $L_{\nu} {\bf 1}(x)=0$ for all $x \in \R^d.$\\
  The unitary equivalence comes from the ground state transformation
 $L^2(\mathbb{R}^d;\nu)\rightarrow L^2(\mathbb{R}^d)$ given by $f \in
 L^2 (\mathbb{R}^d;\nu)\rightarrow f\psi \in L^2 (\mathbb{R}^d),$
 which on dense domain maps $-L_{\nu}$ into $H=\psi{(-L_{\nu})} \,\psi^{-1}$ and yields, 
 by computation for smooth $f$:
 \begin{equation}\label{HAD}
-\psi^{-1}H\,\psi \,f=L_{\nu}\,f=(\triangle +\beta_{\psi}\cdot
 \nabla)f\,,
 \end{equation}
 where we used the expression for $L_{\nu}$ on smooth functions.\\
From (\ref{HAD}) we deduce, multiplying both sides by the smooth function $\psi$, for $f$ smooth:
\begin{equation}\label{HID}
-H\,(\psi \,f)=\psi(\triangle +\beta_{\psi}\cdot
 \nabla)f\,.
 \end{equation}
We claim that, on smooth functions,\begin{equation} H=-\triangle+\tilde{V},\end{equation} with 
\begin{equation}\label{ZI}
\tilde{V}(x):= {1 \over 4}\mid\beta\mid^2_{\psi}(x)+{1 \over 2}\,\mbox{div}\,\beta_{\psi}(x),
 \end{equation}
  In fact then using Leibniz formula and the definition of $ \beta_{\psi}$:
 \begin{eqnarray} (-\triangle+\tilde{V} )\,(\psi\,f)&=&-\triangle
 \psi\cdot f-2 {\nabla
 \psi} \nabla f-\psi\,\triangle f +\frac{\mid\nabla \psi\mid^2}{\psi^2}\,\psi f+\,\mbox{div}\,\frac{\nabla \psi}{\psi}\,\psi f
 \nonumber\\&=& -\triangle \psi\,f-2\, {\nabla \psi}\,\nabla f-\psi\triangle f+  \frac{\mid\nabla
  \psi\mid^2}{\psi}\,f+ \frac{\psi\triangle f-\mid\nabla
    \psi\mid^2}{\psi^2}\,\psi f,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,   
 \end{eqnarray}
 which is seen to be equal to the right hand side of (\ref{HID}).  This shows that (\ref{HAD}) holds.\\
 From (\ref{HID}) we see, taking $f= {\bf 1},$ that $ H\psi=0,$ or using  $H=-\triangle +\tilde{V}$, that $\tilde{V} $ is also expressed by $ \tilde{V}= \frac{\triangle \psi}{\psi}.$\\
Computing in our special case of $\psi$ of the form (\ref{TUN}) and using (\ref{ZIT}) and (\ref{ZI})
 we get 
 \begin{equation}
  \tilde{V}(x)={1 \over 4} \mid\beta_{\varphi}(x)- \omega x \mid^2+ {1 \over 2}\,\mbox{div}\,\beta_{\varphi}(x)-\frac{\omega\,d}{2}\,,\,\,(since\,\,\, \mbox{div}\, x =d).
 \end{equation}
 Hence 
 \begin{equation}
   \tilde{V}(x)=V_{\varphi}(x) + {1 \over 4}\,\omega^2 \mid x\mid^2-{\omega \over 2}\beta_{\varphi}(x)\,x -{\omega \over 2}\,d,
  \end{equation}
 where 
  \begin{equation}
  {V}_{\varphi}(x):={1 \over 4} \mid\beta_{\varphi}(x)\mid^2+ {1 \over 2}\,\mbox{div}\,\beta_{\varphi}(x)
  \end{equation}
 is as stated in the theorem.\\
  Then this proves i), except for the statements on the smoothness of
 the density of the kernel of $p_t,$ for $t >0,$ that follow from the
 general theory of strictly elliptic generators of diffusions, with smooth
 coefficients, and the theory of Dirichlet forms.\\
 ii) The results on the asymptotic expansion for $k$ follow from the one on $v$ as given in theorem \ref{PAP}.
 \end{proof}
 \begin{rem}
i) From the proof it follows that we have also $\tilde{V}=\frac{\triangle \psi}{\psi} $ and $ {V}_{\varphi}=\frac{\triangle \varphi}{\varphi}.  $\\
ii) The unitary transformation $L^2(\mathbb{R}^d)\rightarrow
L^2(\mathbb{R}^d, |\psi|^2\,dx)$ given in the proof of Theorem 3.1,
defined by
\begin{equation} 
f\in L^2 (\mathbb{R}^d)\rightarrow \frac{f}{\psi} \in L^2 \, (\mathbb{R}^d,
|\psi|^2\,dx),
\end{equation} is called ground state transformation, see, e.g,
\cite{Al, AHKS, AlMa, RS}. It permits to transform the operator\,
$-(\triangle + \beta \cdot \nabla)$ with $\beta$ a gradient field( as
coefficient of the first order differential term) into the operator  $-\triangle
+\tilde{V},$ (with a potential term $\tilde{V},$ a multiplication operator, i.e. a zero order differential operator). In probability
theory it is often also called Doob's $h$-transform.
\end{rem}
\section{Some remarks on applications}
We can apply the above theorem \ref{PAAP} to stochastic differential equations descriving  a drift term perturbations of  Brownian motion where the drift is a gradient
field of the form\, $\beta(x)=\beta_{(V)}(x)-\omega\,x,\, \omega >0.$ $\beta_{(V)} $ satisfies the smoothness assumption for $V= ({1 \over 4}\mid\beta\mid^{2}+ {1 \over 2}\mbox{div}\, \beta)$
given in hypothesis 2.1 in terms of the measure $\mu_{V}$ of which $V$ is the
Fourier transform. Such models can be looked upon as perturbations of the
Ornstein-Uhlenbeck (velocity) process by a drift term given by $\beta_{(V)}$.
They have numerous applications e.g. in quantum mechanics and
(quantum) statistical mechanics, see, e.g \cite{AHKS}.\\
 For $d=1$ there are applications in mathematical finance(
smooth perturbations by a drift term of the Vasicek model, for the latter model, see,
e.g. \cite{Vas}).\\
Our results gives essentially small time Borel summable expansions in power series for the
transition probabilities in such models, permitting to compute all
terms, with control on remainders.\\
In areas like mathematical finance, (see, e.g. \cite{FoS, PJA, GHLOW, SA, Wafa}),  models  with multiplicative noise also play an extensive role. They are usually stated for real valued processes satisfying SDEs of the form (1) with $d=1$ but with $\sigma$ not constant  (e.g. models of the type of
Black-Scholes, Pearson's or CIR models).\\
Some of these models can also be covered, by an adaptation of methods in this paper, in the sense of getting for them Borel summable asymptotic expansions for their solutions processes. Let us describe the main
idea, leaving for future publications more detailed applications. The
idea consists in a use of time change in the form of a Lamperti
transformation from an equation of the form (1) for $ X_t$, with $\sigma$
constant positive, to an equation for a transformed process $\tilde{X}_t$
satisfying an equation of the form
$d\tilde{X}_t=\tilde{\beta}_{\sigma}\, (\tilde{X}_{t})\,dt+
dB_{t}$ where $B_t$ is a standard Brownian motion and
$\tilde{\beta}_{\sigma}$ is a new drift obtained by using a Lamperti
transforms $\gamma$ defined by $\displaystyle \gamma_{s_0}(s):=
\int^s_{s_0}\, \frac{dy}{\sigma (y)}, \, {s, s_0}\, \in \mathbb{R},\, s_0< s$ (we assumed $\sigma$ to be smooth strictly positive and such that
$\displaystyle \frac{1}{\sigma} \in L^1_{loc}\, (\mathbb{R}), \,
\frac{1}{\sigma} \notin L^1 (\pm\infty),\, \beta'-\beta
\frac{\sigma'}{\sigma}-\frac{1}{2}\,\sigma\sigma''$ bounded,
$|\beta|,\sigma$ linearly bounded.\\
 By results in Lamperti \cite{La}(see also, e.g.,[\cite{KaShre},Ex. 2.20], \cite{MMH}), we then have
$\tilde{\beta}_{\sigma}(x):=\displaystyle
\frac{\beta(s)}{\sigma(s)}-\frac{\sigma'(s)}{2},$ with $s$ and $x$
related by $\gamma_{s_0}(s)=x$ and the transformation $ x \rightarrow  s$ being invertible, under the stated
assumptions.) The transition function $p^X$ for the process $X$ is
then related to the one for $p^{\tilde{X}}$ by $p^X\,
(t,x,\tilde{x})=\sigma(\tilde{s})^{-1}\, p^{\tilde{X}}\,
(t,s,\tilde{s}),$ with
$\tilde{\gamma}_{s_0}(\tilde{s})=\tilde{x},\,\,s,\tilde{s}, x , \tilde{x}$ being
real variables. Provided $\tilde{\beta}_{\sigma}$ satisfies the
assumptions stated for $\beta$ in theorem \ref{PAAP}, we can apply this theorem to $\tilde{X}$ and then obtain a Borel summable small time asymptotic expansion for $p^{\tilde{X}}$,
thus also for the transition probability kernel for the process
$X$. In this way we obtain the following:
\begin{theo} \label{TOUN} Suppose $\sigma$ satisfies the assumption $\sigma>0,\,\sigma$
smooth, ${1 \over \sigma}\in L^{1}_{loc}(\mathbb{R}), \, {1 \over \sigma}\not\in
L^1(\pm \infty),$ moreover $\beta'-\beta
\frac{\sigma'}{\sigma}-\frac{1}{2}\sigma\sigma''$ bounded, and
$|\beta|,\, \left|\frac{\beta}{\sigma}-\frac{\sigma'}{2}\right|$ and $
\sigma$ linearly bounded at infinity, and finally ${1 \over 4}\mid\tilde{\beta}_{\sigma}(x)\mid^2+ {1 \over 2}\,div\, \tilde{\beta}_{\sigma}(x)$ required to be of the form $\displaystyle
V_{\varphi}(x)+\frac{1}{4}\,{\omega}\,x^2-{\omega \over 2}\,{\beta}_{\varphi}(x)x-{d \over 2}\omega  $ for some ${\omega}>0,\, \varphi>0,$ where $\varphi, V_{\varphi},  {\beta}_{\varphi}$ are as in Theorem \ref{PAAP}. Then $p^X$ has a Borel summable small time asymptotic expansion in powers of $t$.
\end{theo}
\begin{rem}
All assumptions in Theorem \ref{TOUN} are e.g if the following conditions are satisfied:\\
i) for $\sigma>0 $ smooth of the form
$\sigma(x)=\sigma_0\, +c_1 x+ o(x)$ for some $\sigma_0 >0$ sufficiently
small and $c_1>0$ as $x \rightarrow 0$,  and $\sigma(x)= c_2\,x^{\alpha}+o(x^{\alpha} ),$ for some $ \alpha <1, c_2 > 0$ as $\mid x\mid
\rightarrow \infty$.\\
ii) In addition to $\beta$ being $\,C^1$ and $ \mid \beta\mid$ linearly bounded at infinity, one requires that  $\beta(x) \sim x^{\beta},\, \beta >1$ as $\mid x\mid \rightarrow 0 $.\\
We postpone further discussions of such conditions and applications to forthcoming work, relating also to previous work by \cite{Lu} on Borel summability in the
framework of Black-Scholes type models.
\end{rem}
\newpage
\noindent{\bf Acknowledgements:} This work was supported by King
Fahd University of Petroleum and Minerals under the project SF181-MATH-476. The authors gratefully acknowledge this support.

\medskip
\vskip8cm
\begin{flushleft}
\footnotesize
{\it Boubaker Smii} \\
Dept. Mathematics, King Fahd University of Petroleum and Minerals, \\
Dhahran 31261, Saudi Arabia\\
\medskip
{\it Sergio Albeverio}  \\
 Dept. Appl. Mathematics, University of Bonn,\\
HCM;  BiBoS, IZKS, KFUPM(Dhahran); CERFIM (Locarno)\\
\medskip

\medskip
 E-mail: boubaker@kfupm.edu.sa  \\
\hspace{1,1cm}albeverio@uni-bonn.de \\

 \end{flushleft}
\end{document}